OpenStudy (anonymous):

Calculate the amount of agarose and 10x Tris-Gycine buffer to make a 2.5% agarose 1x tris-gycine gel at a final volume of 40ml

OpenStudy (anonymous):

0.025*40mL = 1g of agarose create a ratio you want to go from 10x to 1x so just do a 10 fold dilution, thus 4mL into 40mL should dilute it by 10 fold 10/1 = 40/x x = 4mL